Carol, Debbie and Frieda practice the piano for a total of 23 hours per week. Carol practices twice as much as Frieda, and Debbie practices 3 hours more than Frieda every week. How many hours does each woman practice?
Answer by rfer(16322)

c=2x=10 hrs
f=x=5 hrs
d=x+3=8 hrs
2x+x+x+3=23
4x=20
x=5